### v1

@@ 4 Code and Data
The code developed to implement this framework is publicly available in the form of the Python package naz(Normalizing flow Algorithms beyond Zero-variance training), which can be found at https://github.com/AnaryaRay1/naz. In addition to MAF and CNF, additional architectures such as neural spline coupling and spline autoregressive flows (C. Durkan et al., 2019) are provided, with every model capable of both conditional (supervised) and unconditional (unsupervised) density estimation. The MAF is implemented in both a Pytorch+Pyro (A. Paszke et al., 2017; E. Bingham et al., 2019) and Jax+Numpyro (J. Bradbury et al., 2018; D. Phan et al., 2019) backend with functionality provided for converting one to the other. The Jax+Numpyro backend has been found to be more suitable for Bayesian inference while the Pytorch+Pyro one works better for MLE training. The other architectures are only available with the Pytorch+Pyro backend while the CNF additionally requires the torchdyn (M. Poli et al., 2020) package for implementing the neural ode solvers needed by the FFJORD algorithm. Jax+Numpyro backends for these architectures will be available in future releases of naz. The training data used in the demonstrations is obtained from publicly available BBH PopSynth simulations released in zenodo by, details of which are described in the next section. The trained flows, posterior samples, and smaller training sets will soon be available in zenodo for reproducibility.
5 Illustrative Results: population of compact binary mergers
To demonstrate the performance and accuracy of Bayesian NFs, I use population synthesis simulations of merging BBHs formed through the common envelope evolution of isolated stellar binaries (S. S. Bavera et al., 2020). The formation of these systems was modeled using an earlier version of the POSYDON framework, which combines various stages of binary evolution simulated by the rapid population synthesis code COSMIC (K. Breivik et al., 2020) with the detailed stellar structure calculations performed by MESA (B. Paxton et al., 2011, 2013, 2015, 2018, 2019), and is described in  (T. Fragos et al., 2023). Each population in the simulated dataset is characterized by two initial conditions, namely the common envelop ejection efficiency $\alpha$ and the BH natal spin $\chi_{b}$. The simulations were carried out over a grid of initial conditions spanned by $\alpha\in\{0.2,0.5,1.0,2.0,5.0\}$ and $\chi_{b}\in\{0,0.1,0.2,0.5\}$.
@@ 5 Illustrative Results: population of compact binary mergers
NF emulators can facilitate both SBI and feature prediction from these simulations, which are necessary for gaining astrophysical insights from observed CBC catalogs, now equipped with uncertainty quantification and model selection. To obtain the results presented in this section, I chose the population corresponding to $\alpha=1.0,\chi_{b}=0.1$ to be my test set by excluding it from both training, validation, and Bayesian inference. For the training set, I include 10000 BBHs per grind point for the MLE method used to demonstrate feature prediction and heavily downsample to 20 binaries per grid point for Bayesian inference. Similar conclusions will emerge for less downsampling but will take longer run-times and additional hardware resources to converge for the current MAF-based implementation. To demonstrate how Bayesian NFs can mitigate the biases resulting from un-marginalized flow variance, I also train the same flow using the MLE method on 10 different realizations of the smaller (20 BBHs per grid point) and compare the corresponding density estimates with the posterior predictive emulator provided by Bayesian inference.
@@
The default MAF architecture chosen for all illustrative results comprises 16 masked autoregressive layers, each of which contain 3 hidden layers spanned by 150 units. MLE training was carried with a learning rate decay of 0.5 and early stopping implemented by means of a validation set to minimize over-fitting. With batch sizes of 100, training is fairly efficient on NVIDIA-A100 GPUs with 40GB of memory. For Bayesian inference, however, 4 NVIDIA-H100 GPUs with 80GB of HB200 ram were necessary for running 4 HMC chains with 500 tuning and 1500 sampling steps. These requirements are expected to reduce significantly once Bayesian CNFs are implemented in future explorations.
5.1 Feature prediction using MLE flows for rarely synthesized systems
Figure 1: Feature prediction with PopSynth emulators for under-represented regions of parameter space. Each curve represents the effective spin distribution of BBHs in the test set that belong to an unequal mass ratio and high chirp mass sub-population.
While S. Colloms et al. (2025) have extensively discussed SBI using MLE flows, I show here another astrophysical application of PopSynth emulation, which, like SBI, will also benefit from UQ. Say one is interested in predicting effective spin distributions of a sub-population of BBHs formed through CE, astrophysical parameters same as our test set, that have unequal mass ratio ($<0.3$) and high chirp mass ($>6M_{\odot}$). The CE channel struggles to populate this region of the parameter space (M. Zevin et al., 2020). If as few as $O(10^{4})$ binaries are simulated for this grid point, only $O(10)$ systems fall in this region of parameter space, which are not sufficient for accurate feature prediction.
On the other hand, an NF trained on multiple population runs, each with $10^{4}$ binaries corresponding to different grid points can be used to accurately amplify the dataset of interest, even if it is excluded from training, without having to run a larger PopSynth simulation. In Fig. 1, the mentioned spin distribution is shown for the small simulation set, the amplified set obtained using an NF trained on all available simulations except the one being amplified, and a much larger PopSynth rerun ($10^{6}$ binaries) which is treated as the ground truth. It can be seen that the emulator predicts the distribution with reasonable accuracy despite being trained on a collection of highly downsampled populations. In other words, NFs can be used to predict population features of rarely observed systems without having to simulate a very large number of binaries, which can be prohibitive.
However, training on small synthetic populations per grid point introduces uncertainties that need to be marginalized over to obtain accurate feature prediction (even though data amplification improves feature prediction, deviation from the truth is still observed in Fig. 1, which can in principle be rectified by marginalizing over uncertainties). Similarly, for SBI, when downsampling is necessitated for reducing the computational cost of training a flow on thousands of population runs, log-density evaluation will also be uncertain and hence susceptible to biases unless the variance is accurately quantified and marginalized.
5.2 Uncertainty quantification and model comparison using Bayesian flows
@@
Bayesian inference of the flow parameters yields the posterior predictive distribution of the BBH population that comprises the test set, which is compared with the truth and the prior in Fig. 2. It can be seen that the credible interval of the predicted population enclosed the true distribution, with the posterior being significantly more informative than the prior. Also, it can be seen that the posterior predictive distribution is much closer to the truth than individual MLE reruns on different realizations of the thinned dataset, while also outperforming the prior predictive distribution. This is indicative of the fact that for sparse training data, the posterior predictive emulator should be used for both SBI and feature prediction instead of MLEs. I further compute the BIC for this flow architecture, compare it with that of a different (sub-optimal) one trained on a larger dataset, and list them as follows.
Model Architecture Complexity Data-size BIC
Default 16 layers, 3 hidden layers, 150 neurons 746464 20 binaries per grid 4433736
Simpler 7 layers, 3 hidden layers, 300 neurons 489328 100 binaries per grid 3697137
Note that this not a fair model comparison given that the two models are trained on different data realizations, but rather a demonstration of how model comparison can be implemented. A dedicated model comparison study for different architectures given various PopSynth datasets is left as a future exploration.
@@
To optimize the choice of $\sigma_{0}$ and assess the quantified uncertainties, I show in Fig. 3 calibration curves that compare the empirical coverage of flow predictions with the theoretical coverage for varying numbers of equal probability mass bins. For very small values of $\sigma_{0}$ the uncertainties are underestimated, which indicates that the inference is restricted by the prior. For increasing $\sigma_{0}$, the calibration curves improve, and slight over-estimation is observed for very high values of $\sigma_{0}$ and low values of $n_{Q}$. The overestimation does not seem to increase too much with increasing $\sigma_{0}$, which is expected since broad priors are still consistent with accurate inference as long as there is support near the highest likelihood region. The observed overestimation can be interpreted as the sampler having difficulties converging for a heavily downsampled dataset, a highly complex model, and broad priors. For lower model complexity that guarantees scalable and convergent inference even for broad priors, higher values of $\sigma_{0}$ above a certain threshold are expected to have identical calibration curves. As mentioned before, very high values of $\sigma_{0}$ can lead to the walkers getting stuck at local maxima, causing biases, which the calibration curves would identify as underestimation of error.
@@
Nevertheless, for the complex models used in this demonstration, good calibration is observed for $\sigma_{0}=0.25$, which is consistent with the credible intervals shown in Fig. 2. In other words, the estimated credible intervals indeed encompass the true population distribution the expected number of times, when averaged over all regions of the parameter space, justifying this choice of $\sigma_{0}$. Similar well-calibrated Bayesian NF will correctly quantify the uncertainties in PopSynth emulation from sparse datasets, with the posterior predictive emulators expected to enable unbiased SBI and feature prediction for POSYDONv2 datasets, leading to reliable astrophysical insights from future GW catalogs.
Figure 3: Calibration curves for various choices of the prior scale $\sigma_{0}$ and different number of quantile bins
@@ 6 Discussion and Future Prospects
In this work, I developed a self-consistent method for quantifying epistemic and aleatoric uncertainties in the predictions of NF models and for selecting between different flow architectures to reduce overfitting. By relying on Bayesian neural networks to construct flow transformations and the natural log likelihood function assoiated with probabilitic models such as NFs, I have shown that accurate UQ and model comparison is achievable which successfully mitigates the biases resulting from the variance in flow predictions among multiple re-runs with different parameter initializations or an alternative realizations of a sparse and noisy training sets. I have demonstrated this in the context of population synthesis simulations of merging BBHs, whose emulation is necessary for insightful astrophysical inference from observed catalogs. The developed methodology will be necessary for marginalizing over model uncertainties, which are expected to be non-negligible for sparse PopSynth datasets, which might be unavoidable for high fidelity (and hence costly) simulators such as POSYDONv2 or rarely synthesized sub-populations such as extreme mass-ratio binaries with compact objects in the purported lower-mass gap.
@@
I have discussed several planned improvements in terms of scalability that are necessary for fully realizing the potential of Bayesian NFs as a robust and tractable tool for UQ in simulation-based inference from observed astrophysical populations, which include the use of CNFs to reduce complexity (appendix A.1), approximate Bayesian modeling using stochastic variational inference and importance sampling (appendix A.2), and alternative approaches such as NFs with Monte Carlo Dropouts (appendix A.3). I have further alluded to more general applications of Bayesian NFs in UQ for deep predictive and generative modeling in the context of likelihood-free parameter estimation (appendix B.1) with potential applications to other simulation-based approaches in GW astronomy (M. Dax et al., 2021, 2025), implementations of which are also left as future explorations. As we enter the era of large datasets, high fidelity simulators, and powerful deep learning models capable of combining the two into new scientific discoveries, fully realized Bayesian NFs will play a crucial role in mitigating biases arising from un-marginalized model uncertainties and over/under fitting arising from sub-optimal architecture choices, there by facilitating reliable and accurate astrophysical inference and feature prediction.
@@
I am thankful to Vicky Kalogera for her insights and support throughout this project. I thank Aggelos Katsaggelos, Santiago Lopez Tapia, Ryan Mgee, Deep Chatterjee, Tri Nguyen, Kyle Rocha, Ugur Demir, and Philipp Srivastava for valuable discussions and inputs. This work was supported by the National Science Foundation (NSF) award PHY-2207945. I gratefully acknowledge the support of the NSF-Simons AI-Institute for the Sky (SkAI) via grants NSF AST-2421845 and Simons Foundation MPS-AI-00010513. This research was also supported in part through the computational resources and staff contributions provided for the Quest high-performance computing facility at Northwestern University, which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology. This research used the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois. DeltaAI is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications. I am grateful for the computational resources provided by the LIGO laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459.
@@ A.1 MAF vs CNF: complexity and expressivity
As mentioned before, CNF algorithms can estimate densities with DNNs of much smaller complexity than MAFs. In Fig. 4, I show the performance of a much simpler MAF architecture (2 autoregressive layers, each with 4 hidden layers spanned by 150 units) than the default one, along with a CNF (trained using the FFJORD algorithm, Y. Gal & Z. Ghahramani (2016)) which has a multi-layered perceptron composed of just three 64-unit layers for its neural network. The complexity (number of free parameters) of the CNF is hence orders of magnitude smaller than even this simpler MAF. It can be seen how the simpler MAF struggles to estimate the conditional densities, whereas the CNF performs much better. This implies that a Bayesian CNF emulator for PopSynth will involve sampling a posterior of at most $O(10^{4})$ parameters, which is well within the scope of NUTS. Hence, emulating higher-dimensional BBH populations, the use of broad and flexible priors will be feasible and consistent with convergent Bayesian inference, once the Bayesian CNF is implemented into production. Such a development is left as a future exploration.

### v2

@@ 8 Code and Data
The code developed to implement this framework is publicly available in the form of the Python package naz(Normalizing flow Algorithms beyond Zero-variance training), which can be found at https://github.com/AnaryaRay1/naz. In addition to MAF and CNF, additional architectures such as neural spline coupling and spline autoregressive flows (C. Durkan et al., 2019) are provided, with every model capable of both conditional (supervised) and unconditional (unsupervised) density estimation. The MAF is implemented in both a Pytorch+Pyro (A. Paszke et al., 2017; E. Bingham et al., 2019) and Jax+Numpyro (J. Bradbury et al., 2018; D. Phan et al., 2019) backend with functionality provided for converting one to the other. The Jax+Numpyro backend has been found to be more suitable for Bayesian inference, while the Pytorch+Pyro one works better for standard (MLE) training. The other architectures are only available with the Pytorch+Pyro backend while the CNF additionally requires the torchdyn (M. Poli et al., 2020) package for implementing the neural ode solvers needed by the FFJORD algorithm. Jax+Numpyro backends for these architectures will be available in future releases of naz. The training data used in the demonstrations is obtained from publicly available BBH PopSynth simulations released in zenodo by S. S. Bavera et al. (2020), details of which are described in the next section. The trained flows, posterior samples, and smaller training sets will soon be available in zenodo for reproducibility.
9 Illustrative Results
@@
NF emulators can facilitate both SBI and feature prediction from simulation sets described in Sec. 4, which are necessary for gaining astrophysical insights from observed CBC catalogs, now equipped with uncertainty quantification and model selection. For the training set, we include 10000 BBHs per grind point for the MLE method used to demonstrate feature prediction and heavily downsample to 20 binaries per grid point for Bayesian inference. Similar conclusions will emerge for less downsampling, but will take longer run-times and additional hardware resources to converge for the current MAF-based implementation. To demonstrate how Bayesian NFs can mitigate the biases resulting from un-marginalized flow variance, we also train an ensemble of 150 flows using the MLE method on many different realizations of the smaller (20 BBHs per grid point) and compare the corresponding density estimates with the posterior predictive emulator provided by Bayesian inference. We further show that the Bayesian uncertainty estimates are better calibrated than those obtained through an ensemble of MLE re-runs for such sparse datasets.
@@
The default MAF architecture chosen for all illustrative results comprises 16 masked autoregressive layers, each of which contains 3 hidden layers spanned by 150 units. MLE training was carried with a learning rate decay of 0.5 and early stopping implemented by means of a validation set to minimize over-fitting. With batch sizes of 100, training is fairly efficient on NVIDIA-A100 GPUs with 40GB of memory. For Bayesian inference, however, 4 NVIDIA-H100 GPUs with 80GB of HB200 ram were necessary for running 4 HMC chains with 500 tuning and 1500 sampling steps. These requirements are expected to be reduced significantly once Bayesian CNFs are implemented in future explorations.
We note that these results demonstrate the feasibility of exact posterior sampling for Bayesian normalizing flows, leading to constraints that are more informative than the prior, and show that well-calibrated UQ and potential model comparison between architectures is achievable through this approach. They further illustrate how the quantified uncertainties can mitigate against biases that might result from un-marginalized variance in flow predictions. In addition, they demonstrate the data augmentation capabilities of flow-based PopSynth emulators and how UQ can enable reliable feature prediction in rarely synthesizable sub-populations without having to simulate a prohibitively large number of systems. The fully scalable implementation of this method using Bayesian continuous normalizing flows with robust convergence studies is left as a future exploration.
9.1 Uncertainty quantification in density evaluation using Bayesian flows
@@
To visualize UQ in density evaluation, we emulate the two-dimensional population of BBH chirp masses and effective inspiral spins, since higher-dimensional densities are difficult to represent and marginalize over grids. Bayesian inference of the flow parameters yields the posterior predictive distribution of the BBH population that comprises the test set, which is compared with the truth and the prior in Fig. 2. It can be seen that the credible interval of the predicted population enclosed the true distribution, with the posterior being significantly more informative than the prior. Also, it can be seen that the posterior predictive distribution is much closer to the truth than individual MLE reruns on different realizations of the thinned dataset, while also outperforming the prior predictive distribution. This is indicative of the fact that for sparse training data, the posterior predictive emulator should be used for both SBI and feature prediction instead of MLEs.
9.2 Uncertainty quantification in sample generation using Bayesian flows
Figure 3: Uncertainty quantification for sample generation using Bayesian flows, for emulating the joint distribution of four BBH parameters
For sample generation, it is straightforward to represent and marginalize higher-dimensional densities since the latter amounts to ignoring samples of the marginalized dimensions. Hence, for this demonstration, we emulate the four-dimensional distribution of BBH component masses, effective spins, and redshifts (exactly similar to the emulators of S. Colloms et al. (2025)). As before, Bayesian inference yields posterior samples of the flow parameters corresponding to each of which, we draw samples of BBH parameters and represent their densities as histograms. Figure 3 shows the Bayesian credible intervals on the density of histogram heights in each bin, compared with the truth and the prior. As in the case of density evaluation, it can be seen that the Posterior is encloses the true distribution of the test set while also being considerably more informative than the prior. In other words, the proposed method correctly quantifies uncertainty in sample generation by NF emulators.
9.3 Data augmentation for feature prediction with UQ
As shown in Sec. 5, NF emulators can augment sparse PopSynth datasets to boost sample-statistics without having to simulate a prohibitive number of systems. However, training on small synthetic populations per grid point introduces uncertainties that need to be marginalized over to obtain accurate feature prediction (even though data augmentation improves feature prediction, deviation from the truth is still observed in Fig. 1, which can in principle be rectified by marginalizing over uncertainties). With Bayesian flows, we quantify these uncertainties in data augmentation by predicting a credible interval (an ensemble of augmented datasets) that encloses both the truth and the uncertain prediction from the MLE flow, as can be seen in figure 4.
Figure 4: Feature prediction with uncertainty quantification.
9.4 Model comparison
To assess the efficiency of different flow architectures and avoid overfitting, we compute the BIC for our default architecture and compare it with that of a different (more complex) one, both for the four-dimensional emulator, and list them as follows.
Model Architecture $\log p_{max}(\vec{d}_{train}|\vec{\omega})$ BIC
Default 16 layers, 3 hidden layers, 150 neurons 104 4519833
Simpler 17 layers, 3 hidden layers, 150 neurons -174 9377420
Note that the high complexity of the MAF implies that the BICs are likely driven by the complexity term as compared to the maximum likelihood one. Nevertheless, it can be seen that the model with higher complexity also has a lower maximum log-likelihood achieved during posterior sampling. A dedicated and meaningful model comparison study for different architectures wit Bayesian CNFs, given various PopSynth datasets, is left as a future exploration.
9.5 Calibration of uncertainties
@@
To optimize the choice of $\sigma_{0}$ and assess the quantified uncertainties, we show in Fig. 5 calibration curves that compare the empirical coverage of flow predictions with the theoretical coverage for varying numbers of equal probability mass bins. For very small values of $\sigma_{0}$ the uncertainties are underestimated, which indicates that the inference is restricted by the prior. For increasing $\sigma_{0}$, the calibration curves improve, and slight over-estimation is observed for very high values of $\sigma_{0}$. The overestimation does not seem to increase too much with increasing $\sigma_{0}$, which is expected since broad priors are still consistent with accurate inference as long as there is support near the highest likelihood region. The observed overestimation can be interpreted as the sampler having difficulties converging for a heavily downsampled dataset, a highly complex model, and broad priors. For lower model complexity that guarantees scalable and convergent inference even for broad priors, higher values of $\sigma_{0}$ above a certain threshold are expected to have identical calibration curves. As mentioned before, very high values of $\sigma_{0}$ can also lead to all the walkers getting stuck at local maxima, causing biases, which the calibration curves would identify as underestimation of coverage.
@@
Nevertheless, for the complex models used in this demonstration, good calibration is observed for $\sigma_{0}=0.25$, which is consistent with the credible intervals shown in figures 2, 3. Furthermore, the Bayesian UQ outperforms the ensemble method used by C. Plunkett et al. (2025) which either over or under-estimates coverage for the test set. In other words, the estimated credible intervals indeed encompass the true population distribution the expected number of times, when averaged over all regions of the parameter space, justifying this choice of $\sigma_{0}$. Such well-calibrated Bayesian NFs will correctly quantify the uncertainties in PopSynth emulation from sparse datasets, with the posterior predictive emulators expected to enable unbiased SBI and feature prediction for POSYDONv2 datasets, leading to reliable astrophysical insights from future GW catalogs.
Figure 5: Calibration curves for Bayesian flows (left and center) withvarious choices of the prior scale $\sigma_{0}$ and an ensemble of flows trained using the MLE method (right).
@@ 10 Discussion and Future Prospects
In this work, we developed a self-consistent method for quantifying epistemic and aleatoric uncertainties in the predictions of NF models in the context of PopSynth emulators and potentially for selecting between different flow architectures to reduce overfitting. By relying on Bayesian neural networks to construct flow transformations and the natural log likelihood function associated with probabilitic models such as NFs, we have shown that well-calibrated UQ and plausible model comparison are achievable. Our uncertainty estimates can successfully mitigate the biases resulting from variance in flow predictions among multiple re-runs with different parameter initializations or alternative realizations of a sparse and noisy training sets, while also being better calibrated than the uncertainty estimates obtained from the ensemble. We have demonstrated this in the context of PopSynth simulations of merging BBHs, whose emulation is necessary for insightful astrophysical inference from observed catalogs. We have further shown that NF emulators equipped with UQ can accurately augment sparse PopSynth datasets for feature prediction in rarely synthesizable sub-populations without having to simulate a prohibitively large number of systems.
The developed methodology will be necessary for marginalizing over model uncertainties, which are expected to be non-negligible for sparse PopSynth datasets, which might be unavoidable for high fidelity (and hence costly) simulators such as POSYDONv2 or rarely synthesized sub-populations such as extreme mass-ratio binaries with compact objects in the purported lower-mass gap, and high metallicity binary stars that lead to very few merging BBH systems. In an ongoing investigation, we are exploring the performance of the developed emulators on a diverse collection of POSYDONv2, runs, which will facilitate robust simulation-based probes of binary stellar evolution from observed CBC catalogs, both employing direct SBI or through the guided development and inference of targeted phenomenological population models. Novel astrophysical insights free of biases arising from emulator uncertainties will be plausible.
@@
We have discussed several planned improvements in terms of scalability that are necessary for fully realizing the potential of Bayesian NFs as a robust and tractable tool for UQ in simulation-based inference from observed astrophysical populations. These include the use of CNFs to reduce complexity (appendix A.1), approximate Bayesian modeling using stochastic variational inference and importance sampling (appendix A.2), and alternative approaches such as NFs with Monte Carlo Dropouts (appendix A.3). We have further alluded to more general applications of Bayesian NFs in UQ for deep predictive and generative modeling in the context of likelihood-free parameter estimation (appendix B.1) with potential applications to other simulation-based approaches in GW astronomy (M. Dax et al., 2021, 2025), implementations of which are also left as future explorations. As we enter the era of large datasets, high fidelity simulators, and powerful deep learning models capable of combining the two into new scientific discoveries, fully realized Bayesian CNFs will play a crucial role in mitigating biases arising from un-marginalized model uncertainties and over/under fitting arising from sub-optimal architecture choices, there by facilitating reliable and accurate astrophysical inference and feature prediction.
@@
We are thankful to Vicky Kalogera for her insights and support throughout this project. We are grateful to Mathew Mould for a very detailed, informal review of the manuscript. We thank Aggelos Katsaggelos, Santiago Lopez Tapia, Ryan Mgee, Deep Chatterjee, Tri Nguyen, Kyle Rocha, Ugur Demir, and Philipp Srivastava for valuable discussions and inputs. This work was supported by the National Science Foundation (NSF) award PHY-2207945. We gratefully acknowledge the support of the NSF-Simons AI-Institute for the Sky (SkAI) via grants NSF AST-2421845 and Simons Foundation MPS-AI-00010513. This research was also supported in part through the computational resources and staff contributions provided for the Quest high-performance computing facility at Northwestern University, which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology. This research used the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois. DeltaAI is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications. We are grateful for the computational resources provided by the LIGO laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459.
@@ A.1 MAF vs CNF: complexity and expressivity
As mentioned before, CNF algorithms can estimate densities with DNNs of much smaller complexity than MAFs. In Fig. 6, we show the performance of a much simpler MAF architecture (2 autoregressive layers, each with 4 hidden layers spanned by 150 units) than the default one, along with a CNF (trained using the FFJORD algorithm, Y. Gal & Z. Ghahramani (2016)) which has a multi-layered perceptron composed of just three 64-unit layers for its neural network. The complexity (number of free parameters) of the CNF is hence orders of magnitude smaller than even this simpler MAF. It can be seen how the simpler MAF struggles to estimate the conditional densities, whereas the CNF performs much better. This implies that a Bayesian CNF emulator for PopSynth will involve sampling a posterior of at most $O(10^{4})$ parameters, which is well within the scope of NUTS. Hence, emulating higher-dimensional BBH populations, the use of broad and flexible priors will be feasible and consistent with convergent Bayesian inference, once the Bayesian CNF is implemented into production. Such a development is left as a future exploration.